OpenStudy (lunar-rage):

I need help changing r=3sin theta - 2costheta to rectangular form.

OpenStudy (freckles):

use \[x^2+y^2=r^2 \\ r \cos(\theta) =x \\ r \sin(\theta)=y \\ \frac{y}{x}=\tan(\theta)\]

OpenStudy (freckles):

you only really need 3 of the mentioned

OpenStudy (freckles):

hint: multiply by r on both sides and what you should do next should be more obvious